# AraDiCE: Benchmarks for Dialectal and Cultural Capabilities in LLMs

### v1

@@ 1 Introduction
Large Language Models (LLMs) have consistently pushed the boundaries of Natural Language Processing (NLP), achieving state-of-the-art performance on a wide range of tasks. These models have excelled in areas such as machine translation, summarization, sentiment analysis, and even more complex applications like legal document analysis and creative writing OpenAI (2023); Touvron et al. (2023); Bubeck et al. (2023). Their remarkable ability to extract, reason, and generalize knowledge is fueled by training on vast amounts of data covering diverse topics and domains. However, the success of these models is heavily skewed towards languages with abundant resources, such as English Bang et al. (2023); Ahuja et al. (2023). Low-resource languages, including Arabic and its various dialects, are significantly underrepresented in the datasets used to train these models. This disparity poses a substantial challenge, as LLMs require extensive and diverse data to perform effectively. Consequently, speakers of low-resource languages are at a disadvantage, unable to fully benefit from the advancements in NLP technologies. Recent efforts have been made to train Arabic LLMs Sengupta et al. (2023) and to adapt multilingual models for Arabic Huang et al. (2024); Touvron et al. (2023), but these models are predominantly tailored to MSA, leaving them less effective in handling dialectal Arabic (DA). To formalize this observation, we conduct a systematic study to benchmark Arabic and multilingual models in their performance on Dialectal Arabic. More specifically, we aim to answer the following questions:
@@
To this end, we compiled a comprehensive suite of both existing and newly developed benchmarks to assess the capabilities of these models. We primarily evaluate their fundamental NLP abilities in encoding dialects through dialect identification (e.g. El-Haj et al. (2018)) and dialectal machine translation (e.g. Abdul-Mageed et al. (2023)) tasks. To assess broader cognitive abilities, we include benchmarks that gauge World Knowledge (ArabicMMLU Koto et al. (2024)), Common Sense Reasoning (PIQA Bisk et al. (2020), OBQA Mihaylov et al. (2018), Winogrande Sakaguchi et al. (2021)), Reading Comprehension, BoolQ Clark et al. (2019), Belebele Bandarkar et al. (2024)), and the ability to handle Misinformation (TruthfulQA Lin et al. (2022)). These benchmarks are primarily available in English, but researchers and practitioners have increasingly turned to synthetic data creation methods to address the gap in low-resource languages Long et al. (2024). One promising approach involves utilizing MT to generate synthetic datasets for these languages Sengupta et al. (2023). This method harnesses the capabilities of existing translation models to create large-scale, high-quality synthetic data, which can then be used to train LLMs. In this paper, we introduce a comprehensive approach that leverages MT, specifically from English to MSA and MSA to dialects, combined with human post-editing, to develop synthetic benchmarks for low-resource DA. We concentrate on the Levantine (LEV), and Egyptian (EGY) dialects. Our primary goal is to curate benchmarks that evaluate the performance of LLMs across underrepresented language and dialects. Moreover, we introduce a third and novel benchmark, AraDiCE-Culture, focused on cultural awareness across the Levantine, Egyptian, and Gulf regions. This benchmark evaluates whether LLMs grasp regional cultural nuances beyond language. Our dataset includes questions on public holidays, food, geography, history, public figures, traditional clothing, and more. We probe cultural specifics to assess if the LLMs can differentiate between these regions, emphasizing the importance of models to understand both dialects and their cultural contexts. Our contributions in this work include:
@@
  • We present a comparative analysis of Arabic-focused LLMs, such as Jais Sengupta et al. (2023) and AceGPT Huang et al. (2024), alongside state-of-the-art models like Llama 3 Touvron et al. (2023) and Mistral.
@@
  • We will release the dialectal translation models and benchmarks curated in this study.
@@
LLMs generally struggle with dialect identification, generation, and translation tasks. Although Arabic-centric models like Jais and AceGPT perform better on dialectal tasks, their performance still lags compared to MSA or English. These models often rely on MSA knowledge for distinguishing between dialects. They are better at encoding and understanding dialects than at generating or translating them, as evidenced in dialect generation and translation tasks.
@@
The effectiveness of Arabic-centric models varies significantly across dialects. They perform better in generating Gulf responses, likely due to the dialect’s similarity to MSA. Our findings highlight the importance of AraDiCE in assessing dialectal capabilities and identifying gaps in LLM models.
@@ 2 Datasets
First, we select a number of datasets that have been used for standard LLM benchmarking, including Arabic. The capabilities assessed on these datasets include understanding and generation, world knowledge, commonsense reasoning, reading comprehension, misinformation, and cultural understanding, as listed in Figure 1. We selected the datasets with a focus on diversity and linguistic compatibility. [^2] By linguistic compatibility, we mean that native speakers may not use dialectal Arabic when asking for certain types of information, such as questions related to programming and coding. These datasets include tasks involving generation and multiple-choice questions. The datasets used in this study include: (i) four existing Arabic datasets for understanding and generation: Arabic Dialects Dataset (ADD) El-Haj et al. (2018), ADI Abdelali et al. (2024), and QADI (Abdelali et al., 2021), along with a dialectal response generation dataset Naous et al. (2022), and MADAR Bouamor et al. (2018b); (ii) seven datasets translated into MSA and dialects (Levantine and Egyptian), [^3] Except for ArabicMMLU, which is already in MSA. which include ArabicMMLU Koto et al. (2024) BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), OBQA Mihaylov et al. (2018), Winogrande Sakaguchi et al. (2021), Belebele Bandarkar et al. (2024), and TruthfulQA Lin et al. (2022); and (iii) AraDiCE-Culture, an in-house developed regional Arabic cultural understanding dataset. A detailed description of each dataset is in Appendix A.
@@
We curated a dataset of 180 culturally specific questions by hiring native annotators from the Gulf, Egypt, and the Levant regions to generate seed questions centered on cultural and country-specific themes. After reviewing all submitted questions, we selected those with varying answers across regions. To verify these differences, we appended a country name to each question and used Google Search to examine the top 5 search results. Only questions with distinct answers across countries were retained, resulting in a final set of 30 unique questions. These questions were then translated into the dialects of six countries within the targeted three regions (Gulf, Levant, and Egypt) using GPT-4. The questions span various categories, including public holidays, food, geography, and religion. For gold-standard answers, we followed the NativeQA framework Arid Hasan et al. (2024). Annotators reviewed both the questions and the top 5 search results, combining web data with their cultural knowledge to provide the most accurate answers. Detailed guidelines and sample question-answer pairs are available in Appendices G, A.3 and Table 20.
@@ 3 Machine Translation and Post-Editing
MT is crucial for developing resources for low-resource languages with multiple dialects, like Arabic, which has MSA for formal use and diverse dialects like Egy and Lev for everyday communication. These dialects differ significantly from MSA and from each other, presenting unique challenges for NLP tasks. For MSA, several studies have released datasets, including MMLU, HellaSwag, and Arc-Challenge by Okapi Lai et al. (2023), as well as MMLU by AceGPT Huang et al. (2024). Due to inconsistencies in the number of samples between the original datasets, we refrained from adopting them directly. Instead, we translated these datasets into MSA using Google Translate. [^4] https://pypi.org/project/deep-translator/ Currently, no publicly available translation systems support direct translation between Arabic dialects or between English and specific Arabic dialects; most MT systems focus on MSA-English translation. To address this gap, we trained models to translate from MSA to various Arabic dialects. Translated datasets were manually post-edited for fluency and accuracy. Details of our dialectal MT and post-editing framework are provided below.
@@ 3.1 Dialectal MT
We develop MT models translating between MSA and two major dialects: Egy and Lev Arabic. By translating MSA benchmarks like ArabicMMLU, we create resources for evaluating LLMs’ understanding of dialectal Arabic.
@@
For training our translation systems, we utilized several datasets, including the MADAR Bouamor et al. (2018a), UFAL Krubiński et al. (2023), LDC Mubarak (2018), ArzEn-MultiGenre Hamed et al. (2023), and the SADID Abid (2020). Please see Appendix B.1 for more details.
@@
We fine-tuned two robust machine translation models: AraT5 Nagoudi et al. (2022) and NLLB Team et al. (2022). We experimented with several variants of these models, with sizes ranging from 600M to 3.3B parameters. AraT5 is an Arabic language model based on the T5 (Text-to-Text Transfer Transformer) architecture. In our preliminary experiments, we found the NLLB 3.3B model to surpass AraT5 and its smaller variants, post fine-tuning with dialectal data. We carried out ablation studies using different data mixtures on the NLLB 3.3B model. We shortlisted three systems per dialect using BLEU scores as our primary criterion (See Tables 5, 6 in Appendix B.3) and used human evaluation to select the best system for each dialect. Please see Appendix B for details on our exploration of Dialectal MT.
@@ 3.2 Post-Editing of MT (PEMT)
The translation team comprised 32 native speakers fluent in Lev, Egy, and MSA, with educational backgrounds ranging from Bachelor’s to Master’s degrees, and ages between 21 and 53. Many were professional translators. They received specific guidelines and training tailored to each dataset and dialect to ensure translation quality. To manage the post-editing workload efficiently, we assigned one translator per item, which helped minimize costs and time. A random sample of post-edited texts was reviewed by the expert translators for quality assurance. The process was streamlined using an in-house annotation platform and 17 dedicated projects, resulting in the post-editing of $\approx$45K items. A third-party company handled the hiring and competitive compensation of translators.
@@ 4 Experimental Setup
In our experiments, we employed zero-shot learning to observe performance differences across dialects. Since few-shot learning is known to enhance performance, we did not conduct experiments with it in order to simplify the experimental setups and reduce computational costs. We used prompts in English, MSA, and dialects depending on the task considered and will release both the prompts and configurations publicly.
@@
We used the LM Evaluation Harness [^7] https://github.com/EleutherAI/lm-evaluation-harness for both generation and multiple-choice tasks, employing standard metrics for each task and dataset. We used F1 scores for dialect identification, normalized accuracy for cognitive tasks,888as used in LLM leaderboards https://huggingface.co/docs/leaderboards/open_llm_leaderboard/ and SacreBLEU Post (2018) for machine translation.
@@ 5.2.4 Misinformation
The results for the TruthfulQA task across Lev, Egy, and MSA show that model performances are often close to random, especially for Lev and Egy. Mistral performs best overall, particularly in Lev, but its scores are only slightly above random. We leave a detailed exploration of this for the future.
@@ 5.3 Cultural Understanding
The results on the cultural understanding task (MCQ version) (Figure 9), indicate that Jais is the most culturally aligned model, followed closely by AceGPT, both showing superior awareness of the Egyptian culture. One possible justification is that the Egyptian population (and consequently its online presence) is significantly larger than the other two regions. Llama 3 and Mistral generally show performances close to the random baseline, suggesting limited awareness of Arabic culture. In a qualitative analysis of the models’ responses (when prompting the models in a generation setup), Llama 3 frequently generated fictional entities (e.g., names of people or holidays) and lacked geographical and historical knowledge. For example, when asked about the highest mountain in Jordan, it incorrectly named a mountain in Saudi Arabia. Jais and AceGPT performed better overall, though AceGPT had some issues following instructions as accurately as Jais. Further discussion and example models’ responses can be found in the Appendix E.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have consistently pushed the boundaries of Natural Language Processing (NLP), achieving state-of-the-art performance on a wide range of tasks. These models have excelled in areas such as machine translation, summarization, sentiment analysis, and even more complex applications like legal document analysis and creative writing OpenAI (2023); Touvron et al. (2023); Bubeck et al. (2023). Their remarkable ability to extract, reason, and generalize knowledge is fueled by training on vast amounts of data covering diverse topics and domains.
However, the success of these models is heavily skewed towards languages with abundant resources, such as English Bang et al. (2023); Ahuja et al. (2023). Low-resource languages, including Arabic and its various dialects, are significantly underrepresented in the datasets used to train these models. This disparity poses a substantial challenge, as LLMs require extensive and diverse data to perform effectively. Consequently, speakers of low-resource languages are at a disadvantage, unable to fully benefit from the advancements in NLP technologies. Recent efforts have been made to train Arabic LLMs Sengupta et al. (2023); Bari et al. (2024); Fanar-Team (2024) and to adapt multilingual models for Arabic Touvron et al. (2023); Huang et al. (2024), but these models are predominantly tailored to MSA, leaving them less effective in handling dialectal Arabic (DA). To formalize this observation, we conduct a systematic study to benchmark Arabic and multilingual models in their performance on Dialectal Arabic. More specifically, we aim to answer the following questions:
@@
To this end, we compiled a comprehensive suite of both existing and newly developed benchmarks to assess the capabilities of these models. We primarily evaluate their fundamental NLP abilities in encoding dialects through dialect identification (e.g. El-Haj et al. (2018)) and dialectal machine translation (e.g. Abdul-Mageed et al. (2023)) tasks. To assess broader cognitive abilities, we include benchmarks that gauge World Knowledge (ArabicMMLU Koto et al. (2024)), Common Sense Reasoning (PIQA Bisk et al. (2020), OBQA Mihaylov et al. (2018), Winogrande Sakaguchi et al. (2021)), Reading Comprehension, BoolQ Clark et al. (2019), Belebele Bandarkar et al. (2024)), and the ability to handle Misinformation (TruthfulQA Lin et al. (2022)). These benchmarks are primarily available in English, but researchers and practitioners have increasingly turned to synthetic data creation methods to address the gap in low-resource languages Long et al. (2024). One promising approach involves utilizing MT to generate synthetic datasets for these languages. This method harnesses the capabilities of existing translation models to create large-scale, high-quality synthetic data, which can then be used to train LLMs. In this paper, we introduce a comprehensive approach that leverages MT, specifically from English to MSA and MSA to dialects, combined with human post-editing, to develop synthetic benchmarks for low-resource DA. We concentrate on the Levantine (LEV), and Egyptian (EGY) dialects. Our primary goal is to curate benchmarks that evaluate the performance of LLMs across underrepresented language and dialects. Moreover, we introduce a third and novel benchmark, AraDiCE-Culture, focused on cultural awareness across the Levantine, Egyptian, and Gulf regions. This benchmark evaluates whether LLMs grasp regional cultural nuances beyond language. Our dataset includes questions on public holidays, food, geography, history, public figures, traditional clothing, and more. We probe cultural specifics to assess if the LLMs can differentiate between these regions, emphasizing the importance of models to understand both dialects and their cultural [^2] We follow the definition of culture from AlKhamissi et al. (2024), which describes it as “a multi-faceted inquiry reflecting diversity across worldviews and belief systems.” contexts. Our contributions in this work include:
@@
  • We present a comparative analysis of Arabic-focused LLMs, such as Jais Sengupta et al. (2023) and AceGPT Huang et al. (2024), alongside state-of-the-art models Llama 3 Touvron et al. (2023) and Mistral Jiang et al. (2023).
@@
  • We have released the dialectal translation models and benchmarks curated in this study.
@@
LLMs generally struggle with dialect identification, generation, and translation tasks. Although Arabic-centric models perform better on dialectal tasks, their performance still lags compared to MSA or English. These models often rely on MSA knowledge for distinguishing between dialects. They are better at encoding and understanding dialects than at generating or translating them, as evidenced in dialect generation and translation tasks.
@@
The effectiveness of Arabic-centric models varies notably across dialects, with stronger performance in generating Gulf responses, likely due to the dialect’s closer resemblance to MSA. Our findings highlight the importance of AraDiCE in assessing dialectal capabilities and identifying gaps in LLM models.
@@ 2 Datasets
We begin by selecting a range of datasets commonly used for standard LLM benchmarking, including specific to Arabic. These datasets evaluate capabilities such as understanding and generation, world knowledge, commonsense reasoning, reading comprehension, misinformation, and cultural understanding, as outlined in Figure 1. We selected the datasets with a focus on diversity and linguistic compatibility. [^3] By linguistic compatibility, we mean native speakers may not use dialectal Arabic for queries related to programming and coding. These datasets include tasks involving generation and multiple-choice questions. The datasets used in this study include: (i) four existing Arabic datasets for understanding and generation: Arabic Dialects Dataset (ADD) El-Haj et al. (2018), ADI Abdelali et al. (2024), and QADI (Abdelali et al., 2021), along with a dialectal response generation dataset Naous et al. (2022), and MADAR Bouamor et al. (2018); (ii) seven datasets translated into MSA and dialects (Levantine and Egyptian), [^4] Except for ArabicMMLU, which is already in MSA. which include ArabicMMLU Koto et al. (2024) BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), OBQA Mihaylov et al. (2018), Winogrande Sakaguchi et al. (2021), Belebele Bandarkar et al. (2024), and TruthfulQA Lin et al. (2022); and (iii) AraDiCE-Culture, an in-house developed regional Arabic cultural understanding dataset. A detailed description of each dataset can be found in Appendix A.
@@
We curated a dataset of 180 culturally specific questions by hiring native annotators from the Gulf, Egypt, and the Levant regions to generate seed questions centered on cultural and country-specific themes. After reviewing all submitted questions, we selected those with varying answers across regions. To verify these differences, we appended a country name to each question and used Google Search to examine the top 5 search results. Only questions with distinct answers across countries were retained, resulting in a final set of 30 unique questions. These questions were then translated into the dialects of six countries within the targeted three regions (Gulf, Levant, and Egypt). The questions span various categories, including public holidays, food, geography, and religion. For gold-standard answers, we followed the NativeQA framework Hasan et al. (2024). Annotators reviewed both the questions and the top 5 search results, combining web data with their cultural knowledge to provide the most accurate answers. Detailed guidelines and sample question-answer pairs can be found in Appendices A.3, G and Table 26.
@@ 3 Machine Translation and Post-Editing
Machine Translation is crucial for developing resources for low-resource languages with multiple dialects, like Arabic, which has MSA for formal use and diverse dialects like Egyptian and Levantine for everyday communication Durrani et al. (2014). These dialects differ significantly from MSA and from each other, presenting unique challenges for NLP tasks. For MSA, several studies have released datasets, including MMLU, HellaSwag, and Arc-Challenge by Okapi Lai et al. (2023), as well as MMLU by AceGPT Huang et al. (2024). Due to inconsistencies in the number of samples between the original datasets, we refrained from adopting them directly. Instead, we translated these datasets into MSA using Google Translate. Currently, no publicly available translation systems support direct translation between Arabic dialects or between English and specific Arabic dialects; most MT systems focus on MSA-English translation. To address this gap, we trained models to translate from MSA to various Arabic dialects. Translated datasets were manually post-edited for fluency and accuracy. Details of our dialectal MT and post-editing framework are provided below.
@@ 3.1 Dialectal MT
We develop MT models translating between MSA and two major dialects: Egyptian and Levantine Arabic. By translating MSA benchmarks like ArabicMMLU, we create resources for evaluating LLMs’ understanding of dialectal Arabic.
@@
For training our translation systems, we utilized several datasets, including the MADAR Bouamor et al. (2018), UFAL Krubiński et al. (2023), LDC Mubarak (2018), ArzEn-MultiGenre Hamed et al. (2023), and the SADID Abid (2020). Please see Appendix B.1 for more details.
@@
We fine-tuned two robust machine translation models: AraT5 Nagoudi et al. (2022) and NLLB Team et al. (2022). We experimented with several variants of these models, with sizes ranging from 600M to 3.3B parameters. AraT5 is an Arabic language model based on the T5 (Text-to-Text Transfer Transformer) architecture. In our preliminary experiments, we found the NLLB 3.3B model to surpass AraT5 and its smaller variants, post fine-tuning with dialectal data. We carried out ablation studies using different data mixtures on the NLLB 3.3B model. We shortlisted three systems per dialect using BLEU scores as the primary criterion (see Tables 5 and 6 in Appendix B.3) and conducted human evaluation to select the best system for each dialect. Please see Appendix B for details on our exploration of Dialectal MT.
@@ 3.2 Post-Editing of MT (PEMT)
The translation team comprised 32 native speakers fluent in Levantine, Egyptian, and MSA, with educational backgrounds ranging from Bachelor’s to Master’s degrees, and ages between 21 and 53. Many were professional translators. They received specific guidelines and training tailored to each dataset and dialect to ensure translation quality. To manage the post-editing workload efficiently, we assigned one translator per item, which helped minimize costs and time. A random sample of post-edited texts was reviewed by the expert translators for quality assurance. The process was streamlined using an in-house annotation platform and 17 dedicated projects, resulting in the post-editing of $\approx$45K items. A third-party company handled the hiring and competitive compensation of translators. While synthetic datasets, such as those generated through machine translation, offer valuable opportunities to expand data coverage, they can introduce inherent biases that are difficult to completely mitigate during post-editing. To address this, our approach emphasizes refining translations to ensure alignment with Arabic linguistic nuances and cultural relevance. For tasks requiring world knowledge, we opted for ArabicMMLU (originally in Modern Standard Arabic, or MSA) instead of MMLU, as it better represents the knowledge and contexts relevant to Arabic-speaking communities and regions. For the Winogrande dataset, we made specific adjustments to culturally sensitive instances, as detailed in Section H.1.3.
@@ 4 Experimental Setup
In our experiments, we employed zero-shot learning to observe performance differences across dialects. Although few-shot learning is known to enhance performance, we chose not to include it to simplify the experiments and minimize computational costs. We used prompts in English, MSA, and dialects depending on the task, and released both the prompts and configurations via lm-harness.
@@
We used the LM Evaluation Harness [^7] https://github.com/EleutherAI/lm-evaluation-harness for both generation and multiple-choice tasks, employing standard metrics for each task and dataset. We used F1 scores for dialect identification, normalized accuracy for cognitive tasks, and SacreBLEU Post (2018) for machine translation.
@@ 5.2.4 Misinformation
The results for the TruthfulQA task across Lev, Egy, and MSA show that model performances are often close to random, especially for Lev and Egy, as reported in Figure 7. Mistral performs best overall, particularly in Lev, but its scores are only slightly above random. We leave a detailed exploration of this for the future.
@@ 5.3 Cultural Understanding
The results on the cultural understanding task (MCQ version–Figure 9), indicate that Jais is the most culturally aligned model, followed closely by AceGPT, both showing superior awareness of the Egyptian culture. One possible justification is that the Egyptian population (and consequently its online presence) is significantly larger than the other two regions. Llama 3 and Mistral generally show performances close to the random baseline, suggesting limited awareness of Arabic culture. In a qualitative analysis of the models’ responses (when prompting the models in a generation setup), Llama 3 frequently generated fictional entities (e.g., names of people or holidays) and lacked geographical and historical knowledge. For example, when asked about the highest mountain in Jordan, it incorrectly named a mountain in Saudi Arabia. Jais and AceGPT performed better overall, though AceGPT had some issues following instructions as accurately as Jais. Further discussion and example models’ responses can be found in the Appendix E.
5.4 Additional Model Evaluations
To validate the applicability of the proposed AraDiCE benchmark, we expand our evaluation to cutting-edge models in Arabic and multilingual language modeling, including Fanar Fanar-Team (2024), Gemma-2-9B Team et al. (2024), Aya-Expanse-8B Üstün et al. (2024), Qwen2.5-7B Yang et al. (2024), Llama-3.1-8B-Instruct, and AceGPT-v2-8B-Chat. These models represent diverse architectures and training paradigms, offering valuable insights into handling complex linguistic phenomena. Fanar, built on Gemma, highlights the potential of Arabic-centric pretraining, while Qwen and Aya reflect innovations in multilingual fine-tuning. Due to resource constraints, we focus on Arabic MMLU and PIQA datasets, which assess performance on diverse, challenging tasks. This extended evaluation underscores AraDiCE’s role in advancing Arabic LLM benchmarks and guiding future developments.
Figure 10: Performance of various models on the Arabic MMLU across MSA, Levantine, and Egyptian.
Figure 11: Performance of various models on the Arabic PIQA across MSA, Levantine, and Egyptian.
ArabicMMLU:
As we observed earlier, the models generally perform better on MSA than on the dialects, reflecting the resource-rich training available for MSA data. Among the models (as shown in Figure 10), Fanar 8.7B stands out as the top performer, showcasing strong capabilities in both MSA and dialects. Its superior performance in MSA can be attributed to its MSA-specific training and its foundation on Gemma2 9B, which provides a robust baseline. Notably, Gemma2 outperforms other Arabic-centric models on dialects in the ArabicMMLU, highlighting its strength in handling dialectal variations.
PIQA:
As shown in Figure 11, in PIQA, a task focused on common-sense reasoning, Fanar 8.7B again outperforms the other models, particularly in Egyptian Arabic. Jais 13B follows closely behind, demonstrating solid generalization across MSA and dialects, though it still lags behind Fanar in dialectal contexts. The Qwen2.5 7B model, while less competitive in MMLU, performs better on PIQA for Levantine and Egyptian, suggesting that its reasoning capabilities are more adaptable to dialectal inputs. Gemma2 9B and Aya 8B show more consistent, though lower, performance across both dialects and MSA. The gap between Gemma2 and Aya is relatively small, with Aya performing slightly better in Egyptian Arabic in both MMLU and PIQA tasks, suggesting that Aya might have some potential in specialized domains. These results highlight the ongoing challenges of Arabic dialectal processing and reinforce the efficacy of AraDiCE as a benchmark that can expose performance gaps across dialects. Moreover, the models tested here exhibit varying degrees of sensitivity to dialectal variations, indicating that future model development and fine-tuning efforts need to take into account dialectal diversity to improve overall Arabic LLMs capabilities.